OpenStudy (anonymous):

In a solution at 25°C, the [H+] is 3.5 × 10−6 M. What is the [OH−]?

OpenStudy (anonymous):

3.5 × 10−20 M 2.9 × 10−9 M 1.0 × 10−7 M 3.5 × 10−6 M

OpenStudy (aaronq):

at 25 degrees celsius, \([H^+]*[OH^-]=1.0*10^{-14}\)

OpenStudy (aaronq):

just rearrange that to solve for \([OH^-]\)

OpenStudy (anonymous):

How do I rearrange it?

OpenStudy (aaronq):

it's algebra, a*b=c, if we solve for b, we have divide both sides by a, \(\dfrac{a*b}{a}=\dfrac{c}{a}\) a's cancel on the left, \(\dfrac{\cancel a*b}{\cancel a}=\dfrac{c}{a}\) we're left with: \(\dfrac{b}{a}=\dfrac{c}{a}\)

OpenStudy (aaronq):

sorry i made a copy and paste error, were left with \(\large b=\dfrac{c}{a}\)

OpenStudy (aaronq):

so for your equation, \( [H^+]∗[OH^−]=1.0∗10^{−14}\) divide both sides by \([H^+]\): \( \dfrac{[H^+]∗[OH−]}{[H^+]}=\dfrac{1.0∗10^{−14}}{[H^+]}\) \( \dfrac{\cancel {[H+]}∗[OH^−]}{\cancel {[H^+]}}=\dfrac{1.0∗10^{−14}}{[H^+]}\) \( [OH^−]=\dfrac{1.0∗10^{−14}}{[H^+]}\)

OpenStudy (aaronq):

if you see all these squares with question marks, refresh the page

OpenStudy (anonymous):

Would I plug 3.5 × 10−6 M for H+?

OpenStudy (aaronq):

yep !

OpenStudy (anonymous):

Okay so bear with me [H+] x [OH-] = 1.0 x 10^-14 [ 3.5 × 10^−6] x [?] = 1.0 x 10^-14 [ 3.5 × 10^−6]: [ 3.5 × 10^−6] x [?] 1.0 x 10^-14 ------------------- = ------------ [ 3.5 × 10^−6] [ 3.5 × 10^−6] [OH-] = 28571428.5 I don't understand

OpenStudy (aaronq):

so you got this at the end, right? 1.0 x 10^-14 [OH^-] = -------------- [ 3.5 × 10^−6]

OpenStudy (anonymous):

Yeah?

OpenStudy (aaronq):

okay, that's good. I think you messed up in writing the numbers in your calculator, because the answer is the same. you wrote \(\huge [OH^-]=\dfrac{1.0 *10^{14}}{ 3.5 *10^{−6}}\) when it should've been \(\huge [OH^-]=\dfrac{1.0 *10^{\color{red}{-14}}}{ 3.5 *10^{−6}}\)

OpenStudy (aaronq):

the answer is \(\large 2.857*10^{-9}\) but you got \(\large 2.857*10^{9}\)

OpenStudy (aaronq):

or something like that

OpenStudy (anonymous):

Thanks! I was really lost

OpenStudy (aaronq):

no problem ! glad i could help ya out